OpenStudy (anonymous):

Solve for x: |3x + 2| = 14

OpenStudy (amoodarya):

|3x + 2| = 14 so 3x + 2 = 14 and 3x + 2 = -14 can you go on ?

OpenStudy (amoodarya):

x=4 ,-16/3

OpenStudy (anonymous):

i i cant

OpenStudy (anonymous):

thats the final answer?

OpenStudy (amoodarya):

yes

OpenStudy (anonymous):

ok thxx